1. What is the frequency of microwaves with wavelength of 35 m?
  2. What is the frequency of infrared with wavelength of 7.5 x 10-9 m?
1
Expert's answer
2021-02-11T04:48:05-0500

1. frequency = c/"\lambda"

frequency = 3×108 /35

frequency = 8.5×106

2. frequency = c/"\lambda"

frequency = 3×108 / 7.5×10-9

frequency = 4×1016